# Inheritune: Training Smaller Yet More Attentive Language Models

### v2

@@ 3 Inheritune: Our Proposed Training Recipe
This section offers a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
Recall that we have established the issue of attention degeneration with two motivating examples, highlighting specific inefficiencies in pre-trained LLMs. In this paper, we turn this challenge into an opportunity to create smaller base LMs that are equally performant, achieving similar or lower validation loss compared to their larger, less efficient counterparts, which we refer to as reference models. Our proposed solution builds on two key insights: (1) the early layers of deep LLMs provide effective model initialization, and (2) multiple lazy layers can be collapsed into fewer layers and re-trained to improve the model capacity.
Setup:
We split the dataset into a training set $\mathcal{D}_{\text{train}}$ and a validation subset $\mathcal{D}_{\text{val}}$. Next, we assume that there exists a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps. We want to train a smaller model $\mathcal{M}_{\text{tgt}}$ with the same or better validation loss (lower is better) compared to its larger counterpart $\mathcal{M}_{\text{ref}}$.
@@
    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $n=k/2$ layers of $\mathcal{M}_{\text{ref}}$, including weights, prediction head, and token embeddings.
@@
    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size and repeat steps 1-2 until the desired performance is achieved.
With our method now formally described, we turn to empirical validation. In the following sections, we present comprehensive results demonstrating $\mathsf{Inheritune}$’s effectiveness across various scenarios, including different model sizes and data regimes. In addition, we conducted an in-depth ablation study to analyze the impact of initialization on performance, providing information on the adaptability of our approach.
@@ 4 Experiments
We evaluated $\mathsf{Inheritune}$ through a series of comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M) and GPT-2 medium (355M) models, Radford et al. (2019) pre-trained on the 9B tokens OpenWebText dataset (Gokaslan & Cohen, 2019). These models are trained with data repetition, meaning data is randomly sampled with replacement during batch creation. This experimental setup is adapted from Liu et al. (2023); Sanyal et al. (2024). For evaluation, we compare model(s) trained with $\mathsf{Inheritune}$ with baselines from three key settings: a) baseline models trained from scratch with random initialization, b) baseline models trained using various zero-shot initialization techniques, and c) baseline models trained with knowledge distillation. Table 10 provides detailed specifications for all models used in our experiments. Finally, we conduct a thorough ablation study of our initialization strategy, focusing on the 16-layer GPT-2 medium variants.
We provide experimental details of our $\mathsf{Inheritune}$ training recipe using a GPT-2 large model as an example; a similar procedure was applied to train other models. Our methodology for applying $\mathsf{Inheritune}$involves the following steps:
@@
    Reference Model: We train the full 36-layer GPT-2 Large model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$. This establishes our benchmark validation loss.
@@
    Model initialization We initialize an 18-layer model ($n=k/2$) using the trained 36-layer model as a reference.
@@
    Training and Evaluation: We train the 18-layer model on $\mathcal{D}_{\text{train}}$ for $T$ steps and evaluate its validation loss.
@@
    Iterative Refinement: If the smaller model’s performance is inferior, we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
Models Layers Initialization Steps Pre-train Downstream (0-shot)
Val loss ($\downarrow$) Wikitext ($\downarrow$) Lambada
GPT-2 Medium 24 rand init 100K 2.81 31.93 36.54
16 rand init 100K 2.86 33.67 34.60
16 rand init 200K 2.83
12 Ours 100K 2.87
14 Ours 100K 2.84
Final Model $\longrightarrow$ 16 Ours 100K 2.81 32.04 35.96
GPT-2 Large 36 rand init 100K 2.85 34.84 34.14
18 rand init 100K 2.97 37.63 30.97
18 rand init 200K 2.84
18 Ours 100K 2.80 35.38 34.64
GPT-2 xLarge 48 rand init 100K 2.65 25.45 39.90
24 rand init 100K 2.69 28.32 38.46
24 rand init 200K 2.62
24 Ours 100K 2.64 25.52 43.30
Table 1: Inheritune achieves superior performance with reduced model size. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, 16-layer GPT-2 Medium) against full-sized counterparts and extended training baselines. Metrics include pre-training validation loss ($\downarrow$), zero-shot Wikitext ($\downarrow$) and Lambada performance. Note: GPT-2 xLarge and Large took one round of training; GPT-2 Medium took three rounds.
Baselines trained from scratch (rand init.) :
Here, we compare our model derived using $\mathsf{Inheritune}$ to models of similar size trained with various zero-shot model initialization and efficient training techniques such as stacking, hybrid stacking, and half-width initialization. We explain these baseline training recipes using GPT-2 large and its variants as an example and apply the same process to other models.
  1. 1.
    Stacking Gong et al. (2019); J. Reddi et al. (2023) is a model initialization and efficient (stage-wise) training recipe. We trained a 9-layer GPT-2 large variant from scratch for 100K steps, then expanded the model to 18 layers by copying the weights from layers 0-8 to layers 9-17. Finally, we re-trained this new 18-layer GPT-2 large variant, using stacking initialization for an additional 100K steps.
  2. 2.
    Hybrid stacking: Hybrid stacking is stacking that utilizes a large pre-trained reference model for initialization instead of using its own pre-trained weights. We took the weights of layers 0-8 from the reference 36-layer GPT-2 large model and expanded it to an 18-layer model by copying the weights to layers 0-17. We then trained this new 18-layer GPT-2 variant for 100K steps.
  3. 3.
    Half width: We initialized the baseline GPT-2 large variant across the width dimension and preserved the entire depth. We copied the weights of the first half of the attention heads (0-9) and MLPs of the GPT-2 large reference model into the baseline GPT-2 variant with half the width but all layers.
Baselines trained with Knowledge Distillation
As a baseline, we first apply logit-based knowledge distillation (KD) Hinton et al. (2015) to train a 16-layer GPT-2 medium variant (student) initialized randomly. For the second baseline, we use a DistillBERT-style approach Sanh et al. (2019), where the 0-11 layers of the student model are initialized with each alternate block of its teacher, and the remaining 4 blocks are initialized using layers 18, 19, 20, and 21 of the teacher. Both baselines are trained for 50K steps, using a vanilla 24-layer GPT-2 medium model as the teacher (our reference model). Additional results with 12-layer GPT-2 medium variant(s) trained from scratch, using KD baseline and our recipe can be found in Figure 8.
@@
Table 2: Inheritune outperforms baseline zero-shot initialization and efficient training techniques. Comparison of pre-training validation loss for GPT-2 xLarge, GPT-2 Large and GPT-2 Medium variants. $\mathsf{Inheritune}$-derived models consistently achieve lower loss compared to models initialized with stacking, hybrid stacking, and half-width techniques.
@@ 4.1 Results and Discussions
Models trained with Inheritune outperforms much larger models trained from scratch.
@@
We present our main results in Table 1. Our 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 xlarge, 36-layer GPT-2 large, and 24-layer GPT-2 medium achieve comparable or lower validation losses than their full-sized counterparts when trained for the same number of steps (100K steps). Our GPT-2 xlarge and GPT-2 large variants undergo one round of $\mathsf{Inheritune}$ training, while for GPT-2 medium, we perform three rounds of training with 12, 14, and 16 layers. We also evaluated all models on two next-word prediction downstream tasks in a zero-shot setting using the Wikitext Merity et al. (2016) and Lambada Paperno et al. (2016) datasets. The downstream performance of our GPT-2 models derived using the $\mathsf{Inheritune}$ recipe matches their much larger reference models. From the convergence perspective, some prior works have made connections between over-parameterization and faster convergence Bengio et al. (2005); Vaswani et al. (2018). In supplementary Figure 5, we show that the small LMs derived with $\mathsf{Inheritune}$, although smaller compared to their reference models, still converge as fast as their large-size reference models.
Models trained with Inheritune outperform same-sized models trained from scratch.
Table 1 demonstrates that GPT-2 variants trained with $\mathsf{Inheritune}$ outperform their same-sized counterparts trained from scratch, both when trained for the same number of steps and even when trained for double the steps (200K). This result underscores the efficiency of our approach. The only exception is the 24-layer GPT-2 xlarge variant, which surpasses both our model and the full-size model when trained for 200K steps.
Models trained with Inheritune outperform all zero-shot model initialization and efficient training baselines.
In Table 2, we compare GPT-2 xlarge, GPT-2 large, and GPT-2 medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid, and half-width initialization baselines. The half-width baseline performs poorly, revealing the limitations of naive width reduction. Although stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$. In all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy. For a detailed view of the training dynamics across all methods, refer to the training curves in supplementary Figure 6.
@@
Figure 3: A 16-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized model trained with Logit-based distillation baselines. We conducted vanilla KD Hinton et al. (2015) and DistillBERT-style KD Sanh et al. (2019) with teacher initialization, using a 24-layer GPT-2 medium as the teacher for both KD baselines.
@@
In Figure 3, we compare 16-layer GPT-2 medium variants derived using vanilla knowledge distillation Hinton et al. (2015) and DistillBERT-style distillation Sanh et al. (2019), which leverages teacher layers for model initialization, vanilla training from scratch and $\mathsf{Inheritune}$. Our model demonstrates faster convergence and significantly better final generalization after 50K steps. Additional distillation experiments can be found in the supplementary material.
@@ 4.2 Ablations
Table 3: Impact of initializing various sub-modules within a transformer block. We compare the validation loss of a 16-layer GPT-2 medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 medium reference model. All models are trained on the OpenWebText-9B dataset. Key findings: (1) $\mathsf{Inheritune}$ initialization and attention + MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact. A detailed training curve is shown in Figure 7.

### v3

@@ 3 Inheritune: Our Proposed Training Recipe
This section provides a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
@@
    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $l$ out of $L$ layers of $\mathcal{M}_{\text{ref}}$, including prediction head, and token embeddings.
@@
    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size by adding next few contiguous layers and repeat steps 1-2 until desired performance is achieved.
As previously established, we have identified the problem of attention degeneration and its connection to lazy layers, highlighting specific inefficiencies in pre-trained LLMs. In this work, we transform this challenge into an opportunity to create smaller base language models, which we refer to as target models $\mathcal{M}_{\text{tgt}}$, that achieve comparable performance with similar or lower validation loss compared to their larger, less efficient counterparts, which we term reference models $\mathcal{M}_{\text{ref}}$.
Algorithm 1 Inheritune: Training Recipe for Small Language Models
1:Reference model $\mathcal{M}_{\text{ref}}$ with $L$ layers, datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{val}}$, steps $\mathsf{T}$
2:Copy embedding layer and LM head from $\mathcal{M}_{\text{ref}}$ to $\mathcal{M}_{\text{tgt}}$
3:Select $l$ early contiguous layers from $\mathcal{M}_{\text{ref}}$ with high AvgRank
4:Initialize $\mathcal{M}_{\text{tgt}}$ with selected layers between embeddings and LM head
5:Train $\mathcal{M}_{\text{tgt}}$ on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps
6:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
7:     Grow $\mathcal{M}_{\text{tgt}}$ by inheriting additional layers
8:     Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps
9:end while
10:return Optimized model $\mathcal{M}_{\text{tgt}}$
Our proposed solution builds on two key insights: (1) the early layers of deep LLMs contain a higher concentration of potent layers with high AvgRank values, making them suitable for effective model initialization, and (2) lazy layers can be identified, removed, or utilized in smaller numbers, then subsequently re-trained to improve overall model capacity.
Figure 4: Overview of the $\mathsf{Inheritune}$ training recipe using a 24-Layer GPT-2 medium model example. A smaller target model is initialized using early layers from a larger, pre-trained reference model. The target model goes multiple rounds of training while inheriting more early layers until it matches the reference model. The intensity of the red color in layers correlates with $\mathrm{MaxRank}^{(l)}$.
Setup:
We split the dataset into a training set $\mathcal{D}_{\text{train}}$ and a validation subset $\mathcal{D}_{\text{val}}$. Next, we assume that there exists a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $L$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{L-1}\}$ trained with $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps. We want to train a smaller model $\mathcal{M}_{\text{tgt}}$ with the same or better validation loss (lower is better) compared to its larger counterpart $\mathcal{M}_{\text{ref}}$.
With our method now formally described, we turn to empirical validation. In the following sections, we present comprehensive results demonstrating $\mathsf{Inheritune}$’s effectiveness across various scenarios, including different model sizes and data regimes. In addition, we conducted an in-depth ablation study to analyze the impact of initialization on performance, providing insights into the adaptability of our approach.
@@ 4 Experiments
We evaluate $\mathsf{Inheritune}$ through a comprehensive set of experiments using several GPT-2 models: 48-layer GPT-2 xlarge (1.5B), 36-layer GPT-2 large (770M), 32-layer GPT-2 large (668M) and 24-layer GPT-2 medium (355M) models Radford et al. (2019). Table 11 provides detailed specifications of all model configurations used.
We use two training datasets in our experiments: OpenWebText Gokaslan & Cohen (2019) and Fineweb_edu Penedo et al. (2024). Our experimental setup closely follows prior works Liu et al. (2023); Sanyal et al. (2024). For models trained on OpenWebText, we report validation loss (log perplexity) on a held-out set. For those trained on Fineweb-Edu, we report training loss (also log perplexity). In addition, we conduct zero-shot downstream evaluations using the lm-evaluation-harness Gao et al. (2024) across five standard tasks: ARC-easy (ARCE; Clark et al. (2018)), LAMBADA Paperno et al. (2016), SciQ Welbl et al. (2017), HellaSwag Zellers et al. (2019), and PIQA Bisk et al. (2020). We also perform a thorough ablation study of our initialization strategy, focusing on 16-layer GPT-2 medium variant.
We provide experimental details our $\mathsf{Inheritune}$ training recipe using a GPT-2 medium model as an example; similar procedure was applied to train other models. A visualization of the training recipe is presented in Figure 4. Our recipe for applying $\mathsf{Inheritune}$ involves the following steps.
@@
    Reference Model: We train the full 24-layer GPT-2 medium model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$. This establishes our benchmark validation loss.
@@
    Model initialization: We initialize an 12-layer model ($l=L/2$) using the reference model.
@@
    Training and Evaluation: We train the 12-layer model on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$=100K steps and evaluate its validation loss.
@@
    Iterative Refinement: If the smaller model’s performance is inferior to the reference model, then we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
We choose $l=L/2$ as the starting point and increase the model size by two layers in each round across all our experiments, aiming to minimize the number of training rounds. In principle, $\mathsf{Inheritune}$ should generalize to other hyperparameter choices as well.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model, despite having much fewer layers. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge variant, 18-layer GPT-2 Large variant, 16-layer GPT-2 Medium variant) against their full-sized counterparts and same sized variants trained from scratch. All models are trained for 100K steps using OpenWebText data with data repetition.
(a) GPT2-Medium
(b) GPT2-Large
Figure 6: Models derived using $\mathsf{Inheritune}$ converge faster and match the final validation loss and average downstream performance (Perf.) across 5 datasets of the full-sized model, while outperforming same-sized models. (Plots) Validation loss curves for our 16-layer variants versus full-sized models. (Table) Zero-shot downstream performance comparison. All models are trained for 100K steps using FineWeb_edu data without data repetition.
Models Recipe Layers Perf. Medium rand init 24 47.74 rand init 16 46.84 Ours 16 48.08 Large rand init 32 49.44 rand init 16 48.17 Ours 16 49.75
Baselines.
We compare GPT-2 model variants (variants are models with lesser layers compared to their vanilla configuration) trained using $\mathsf{Inheritune}$ against two types of baselines: (1) larger reference models with more layers, and (2) similar-sized models with the same number of layers trained from scratch (randomly initialized). Additionally, we compare $\mathsf{Inheritune}$ against efficient training baselines that leverage pre-trained model initialization. These include stacking Gong et al. (2019); J. Reddi et al. (2023), hybrid stacking (which initializes the model using layers from the reference model), and half-width initialization (where all layers are retained, but the hidden dimension and number of attention heads are halved and initialized from the reference model). Finally we briefly compare $\mathsf{Inheritune}$ with knowledge distillation (Hinton et al., 2015). Detailed descriptions of these baselines are provided in the supplementary material Section B.
@@
Table 1: $\mathsf{Inheritune}$ outperforms baseline zero-shot initialization and efficient training techniques. Comparison of pre-training validation loss for GPT-2 xLarge, GPT-2 Large and GPT-2 Medium variants. $\mathsf{Inheritune}$-derived models consistently achieve lower loss compared to models initialized with stacking, hybrid stacking, and half-width techniques.
@@ 4.1 Results and Discussions
Models trained with $\mathsf{Inheritune}$ outperform much larger models trained from scratch.
@@
We present our main results in Figure 5. The 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 xlarge, 36-layer GPT-2 large, and 24-layer GPT-2 medium, respectively, achieve comparable or lower validation losses than both their full-sized counterparts and same-sized models trained from scratch, when trained for the same number of steps (100K). Our GPT-2 xlarge and GPT-2 large variants require a single round of $\mathsf{Inheritune}$ training, while the GPT-2 medium variant undergoes three rounds with 12-, 14-, and 16-layer configurations. From a convergence perspective, prior work has linked overparameterization to faster convergence Bengio et al. (2005); Vaswani et al. (2018). Interestingly, we find that smaller models derived using $\mathsf{Inheritune}$ converge just as fast as their larger counterparts. A detailed summary of these results is provided in the appendix Table 3.
We conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al. (2023); Biderman et al. (2023) using the high-quality Fineweb_edu dataset. We trained a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch. Next, we trained two 16-layer variants: one derived from GPT-2 large and the other from GPT-2 medium, using their respective reference models following $\mathsf{Inheritune}$. For comparison, we also trained 16-layer baseline models from scratch. All models were trained for 100K steps. We observe thematically consistent results: as shown in Figure 6, the 16-layer variants trained with $\mathsf{Inheritune}$ consistently match the performance of their full-sized counterparts and outperform same-sized baselines, both in terms of training loss and zero-shot downstream evaluation. Downstream results are averaged over five tasks (see table within Figure 6); the complete breakdown is provided in Table 5 placed in appendix. Model configurations and training hyper-parameters are detailed in the supplementary material (refer Section F).
@@
Figure 7: A 16-layer GPT-2 medium variant derived using $\mathsf{Inheritune}$ converges faster and generalizes better than a same-sized model trained with logit-based distillation baselines.
Models trained with Inheritune outperform all zero-shot model initialization baselines.
In Table 1, we compare GPT-2 xlarge, GPT-2 large, and GPT-2 medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid, and half-width initialization baselines. The half-width baseline performs poorly, revealing the limitations of naive width reduction. While stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$. Across all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy. For a detailed view of the training dynamics across all methods, refer to the training curves in supplementary Figure 14.
@@
In Figure 7, we compare 16-layer GPT-2 medium variants derived using vanilla knowledge distillation Hinton et al. (2015) and DistillBERT-style distillation Sanh et al. (2019), which leverages teacher layers for model initialization, vanilla training from scratch and $\mathsf{Inheritune}$. Our model demonstrates faster convergence and significantly better final generalization after 50K training steps. Additional distillation experiments can be found in the supplementary materials.
@@ 4.2 Ablations
Table 2: Impact of initializing various sub-modules within a transformer block. We compare validation loss of a 16-layer GPT-2 medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 medium reference model. All models are trained on the OpenWebText dataset. Key findings: (1) $\mathsf{Inheritune}$ initialization and attention + MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact. A detailed training curve is shown in Figure 15.